Solve the equation. x+8 = 0
Answer #1

23 +8 = 0 3 2°+ (2)3=0 -> (x+2) (x2–2x+4) = 0 :. (x+2) = 0, (x2–2x+4)= 0 n. 2= 2 22–2x+4 = 0 x= -b = √b²-4ac 2a -(-2) ± √(-2